OpenStudy (anonymous):

The square root of the quantity 72 times n to the thirteenth power.

OpenStudy (anonymous):

I have no clue how to do this.

OpenStudy (lucaz):

I'm not sure if I get it.. sqrt(72n)^13 ?

OpenStudy (anonymous):

Yes, that's it

OpenStudy (lucaz):

or sqrt(72)(n)^13

OpenStudy (anonymous):

The first one.

OpenStudy (anonymous):

[sqrt(x)]^13 = x^6.5 = x^6 sqrt(x)

OpenStudy (anonymous):

Here are the answer choices : A. 8n square root 3n^6 B. 3n^6 square root 8n C. 8n^6 square root 3n D. 6n^6 square root 2n

OpenStudy (anonymous):

I have like 10 of these do do =(

OpenStudy (lucaz):

right, I'm trying to solve it, you have to use the proporties of exponents http://www.mathplanet.com/education/algebra-1/exponents-and-exponential-functions/properties-of-exponents

OpenStudy (lucaz):

@satellite73 can you help here?

OpenStudy (triciaal):

sqrt (72n^13) find factors of 72 and n that will give perfect sq rewrite sqrt (9 X 4 X 2 X n^12 X n^1) 3 X 2 X n^6 X sqrt2n